OpenStudy (anonymous):

supposed a item sold for was 75 cent per pound.now it's 81 how much percent did it go up to

OpenStudy (solomonzelman):

Oh, so your first cost is `75 cents per item`. Then, the cost has gone up to `81 cents per item`. You want to find the following: by how many percent did the original price went up

OpenStudy (solomonzelman):

I will give you an example: If my first price was `60 cents per item` and the new price was `90 cents per item` THEN to find the "percentage increase" (or "by how many percent did the original price go up") I would do this: `Step 1:` 90-60=30 So I know the price per item has gone up by 30 cents. `Step 2:` 30 ÷ 60= ½ So I know that the price hs gone up by half of the original price. `Step 3:` How many percent is ½? Converting ½ to percent, we get: (1/2)•100=50 `Answer:` The price has gone up by 50%.

OpenStudy (solomonzelman):

But, if you want to know how many percent it has gone up `TO` then, you will add the precents. Your original price 60 cents (per item) is 100% And it has gone up by additional 50% (as we found out) So in that example, it has gone up to 150% of the original price.

OpenStudy (anonymous):

thanks

OpenStudy (solomonzelman):

If you got any questions you are always welcome t ask:)

OpenStudy (solomonzelman):

And if I happen to be not online or unavailable, you can tag any other user. @Ssjmichaelcraig See how I tagged you? Same way you tag others - put @ in front. good luck!

OpenStudy (anonymous):

thanks so much

OpenStudy (solomonzelman):

You are always welcome!